### v3

@@ 1 Introduction and overview
Some other economists, including John Rae (1834) and John Stuart Mill (1848) (see Appendix C), argued that capital growth might also be explained by a rise in productivity of capital and labor already in place. Ways might found for existing factors to produce more, that is, and so to allow more consumption, or more capital growth, or any mix of the two, without inputs of net saving. Robert Solow (1957) allowed that possibility for “disembodied” growth, where plant and products already existing are repurposed or redeployed in more productive ways.
We test between those two explanations of capital growth, by net saving or by increase in productivity of capital and labor already in place, by comparing net saving to concurrent change in market-value capital in 88 countries. As changes in net saving are expected to be associated with opposite changes in consumption, we also compare change in consumption to concurrent change in capital growth (capital acceleration). All data are drawn from national accounts of those countries as collated on the free website World Inequality Database.
@@
If free growth theory is right, tax policy and other policy to encourage saving over consumption should be reconsidered. These policies include the higher tax on ordinary income than on capital gains, and the double tax on corporate dividends.
@@
2 Net saving and capital growth
@@
${}{\Delta K_{i}=S_{net,i}+Q_{i}\ ,}$ (1)
where $Q_{i}$ will mean the sum of market noise, which may prove positive or negative or zero, plus any part of capital growth explained by concurrent productivity gain as described by Rae and Mill, and by Solow as to disembodied growth. We call this sum of noise and productivity gain “free growth” in that it costs no net saving.
The object of testing is to find effects of $S_{net}$ on concurrent $\Delta K$, and so to help evaluate historic and current teachings as to those effects. We submitted above that most teaching, with exceptions noted as to Rae, Mill, and Solow, and within the warranted rate, predicts net saving to differ from concurrent growth in market-value capital only by market noise which tends to balance out over scale and time. The residual term $Q$ in Eq. (1), in that case, will give the market noise converging to zero. That consensus prediction, which we will challenge, will here be called “thrift theory”; $S_{net}$, in thrift theory, is expected to converge to $\Delta K$ if held within the warranted rate. That is,
${}{E(\Delta K)=S_{net}\quad\text{or equivalently}\quad E(Q)=0,\quad\text{if}\quad\frac{\Sigma S_{net}}{\Sigma K}\leq u\ ,\quad\text{in thrift theory,}}$, (2)
@@
$E(\Delta K)$ and $E(Q)$ here give the expected values of $\Delta K$ and $Q$ respectively. Expected value means predicted average of outcomes over all observations. In this case, that will mean predicted average of yearly observations over all years reported. As secular economic growth has tended to make later stocks and flows larger than earlier ones, we first divide by $K$ (normalize) to avoid overweighting of more recent years in finding that average. Division of Eq. (1) by $K_{i-1}$ gives
${}{\frac{{\Delta K}_{i}}{K_{i-1}}=\frac{S_{net,i}}{K_{i-1}}+\frac{Q_{i}}{K_{i-1}}\ .}$ (3)
The first term in Eq. (3) gives capital growth rate $g(K)$. The second term is a variant of the Keynesian net saving rate $s_{net}$ where capital rather than output becomes the denominator. This flow will here be called “thrift”. It will show as $s(K)$, with the subscript “net” left implicit, and with the understanding that the denominator shows capital at market value, rather than at depreciated cost. The third term in Eq. (3) will be called free growth rate and shown as $q(K)$. Then
$g{(K)}_{i}=\frac{\Delta K_{i}}{K_{i-1}}\ ,\quad s{(K)}_{i}=\frac{S_{net,i}}{K_{i-1}}\ ,\quad\text{and}\quad q{(K)}_{i}=\frac{Q_{i}}{K_{i-1}}\ ,$
@@
${}{g{(K)}_{i}=s{(K)}_{i}+{q(K)}_{i}\ .}$ (4)
By the definition $q(K)=\frac{Q}{K}$, an expected value $E(Q)=0$ implies $E\left(q(K)\right)=0$. Application of Eq. (2) to Eq. (4) now gives
${}{E\left(g(K)\right)=s{(K)}\quad\text{and}\quad E(q(K))=0\ ,\quad\text{under thrift assumptions,}}$, (5)
where “thrift assumptions” are that thrift theory is correct and that the warranted rate is not exceeded. Meanwhile Eq. (4) allows
${}{\Delta g{(K)}_{i}=\Delta s{(K)}_{i}+\Delta q{(K)}_{i}\ ,\quad\text{where}}$ (6)
$\Delta g{(K)}_{i}=g{(K)}_{i}-g(K)_{i-1}\ ,\quad\Delta s(K)_{i}=s(K)_{i}-s(K)_{i-1}\ ,\quad\text{and}\quad\Delta q{(K)}_{i}=q{(K)}_{i}-q{(K)}_{i-1}\ .$
@@
${}{E\left(\Delta g(K)\right)=\Delta s(K)\quad\text{and}\quad E(\Delta q(K))=0\ ,\quad\text{under thrift assumptions.}}$. (7)
The first term in Eq. (6) may be called “capital acceleration”. Division of Eq. (6) by capital acceleration, and rearrangement, gives
${}{\frac{\Delta s{(K)}_{i}}{\Delta g{(K)}_{i}}+\frac{\Delta q{(K)}_{i}}{\Delta g{(K)}_{i}}=1\ .}$ (8)
Define $\theta_{s,i}=\frac{\Delta s{(K)}_{i}}{\Delta g{(K)}_{i}}$ and $\varphi_{s,i}=\frac{\Delta q{(K)}_{i}}{\Delta g{(K)}_{i}}$ to restate Eq. (8) as
@@
$\theta$ and $\varphi$ will be called the “thrift index” and “free growth index” respectively. $\theta_{s}$ and $\varphi_{s}$ give their values as found from data for net saving. Expected values, again, are predicted averages of outcomes. Thus Eq. (11) and thrift theory can be tested by finding average values of $\theta_{s,i}$ and comparing findings to the expected value $\theta=1$. First we find
@@
$\overline{\theta_{s,i}}\cong 1\quad\text{and}\quad\overline{\varphi_{s,i}}\cong 0\ .$
@@
Fig. 1 shows average values of $\theta_{s,i}$ and $\varphi_{s,i}$ for 88 countries, both unweighted and weighted to GDP, over the period 1980-2022. To control distortions brought by small absolute denominators, years were screened out where $|\Delta g(K)|$ was found at less than 0.01 (see Section 10). Results show $\overline{\varphi_{s,i}}\cong 1$ and $\overline{\theta_{s,i}}\cong 0$. These findings appear to refute thrift theory, and to support free growth theory as defined earlier. $Q$, predicted in thrift theory to describe effects of market noise converging to zero, is revealed to include also the effects of productivity gain as described by Rae, Mill and Solow. We will now test thrift theory from a different approach.
Fig. 1: Thrift and free growth indexes derived from net saving (88 countries).
Table 1: Regression of $\Delta g(K)$ to value shown (Screen = 0.01). $H_{0}\ \text{per thrift theory:}\ \theta_{s}=1\ \&\ \varphi_{s}=0$
$\theta_{s}$ $\varphi_{s}$
Regression of $\Delta g(K)$ to value shown ${}^{***}$ ${}^{***}$
@@ 3 Consumption and capital growth
Modern growth theory reasons from the causalities of growth in capital, labor and productivity, as represented, for example, in Cobb-Douglas functions. We prefer to find mathematical possibilities first, then to test among them from the data, and then to infer causalities from results of the tests.
In national accounts and most of the literature, saving equals consumption forgone in that a fixed source of spending is spent on the sum of the two. Replacement saving to offset depreciation costs an equal potential gain in consumption, that is, while net saving costs an equal actual decrease in consumption. We accept these meanings as definitions of saving and net saving. Keynes (1936) expressed this idea in the equation $c+s=1$, where $c$ and $s$ are the ratios of consumption and saving to output. We have been dividing by capital rather than output, and we are interested in net saving after capital replenishment, which thrift theory equates to capital growth. If we notate the consumption/capital ratio as $c(K)$, and recall that $s(K)$ means the ratio of net rather than gross saving to capital, we have
@@
$E(\Delta g(K))=\Delta s(K)=-\Delta c(K)\ ,\quad\text{under thrift assumptions.}$. (12)
We will test this expectation from data for consumption. First we define $z(K)$ implicitly in the equation
$g(K)_{i}=z(K)_{i}-c(K)_{i}\ ,\quad\text{from which}$ (13)
$\Delta g(K)_{i}=\Delta z(K)_{i}-\Delta c(K)_{i}\ ,$ (14)
As with $q(K)$ in earlier equations. $z(K)$ will mean any cause other than consumption restraint in explaining capital growth, and hence will mean free growth as measured from consumption data.
It is accepted that capital, output and consumption can all grow at the same constant rate overall (balanced growth), even when capital growth is equated to net saving from consumption, if saving and consumption grow by turns. $s(K)$, that is, might be raised to give more saving in one period, and then dropped to give more consumption in the next. If some households were saving while others were making up for consumption lost, also, then overall $g(K)$, $s(K)$ and $c(K)$ might grow at the same rate continuously. Thus it is impractical to test thrift theory directly from Eq. (13), which allows balanced growth.
But balanced growth is recognized to be impossible in capital acceleration. If net output captures all of value added, that is, while consumption captures all of value surrendered in taste satisfaction, then the output/consumption rate must rise when capital acceleration is positive, and conversely when negative. Therefore we will test thrift theory from Eq. (14) alone. As before, we first divide Eq. (14) by $\Delta g(K)_{i}$ to get
${}{\frac{\Delta z(K)_{i}}{\Delta g(K)_{i}}+\frac{-\Delta c(K)_{i}}{\Delta g(K)_{i}}=1\ .}$ (15)
Define $\theta_{c,i}=\frac{-\Delta c{(K)}_{i}}{\Delta g(K)_{i}}$ and $\varphi_{c,i}=\frac{\Delta z{(K)}_{i}}{\Delta g(K)_{i}}$ to re-express Eq. (15) as
@@
Table 2: Average $\varphi_{s,i}$ and $\varphi_{c,i}$ in all countries (screen = 0.01). Number of years clearing screen shown in ()
@@
Country Period $\overline{\varphi_{s,i}}$ $\overline{\varphi_{c,i}}$ Armenia 1997 - 2018 (17) 0.95 1.00 Aruba 1997 - 2001 (5) 0.77 2.02 Australia 1962 - 2019 (43) 0.97 1.01 Austria 1997 - 2017 (14) 0.92 1.03 Azerbaijan 1997 - 2018 (20) 0.67 1.08 Bahrain 2010 - 2013 (4) 0.61 1.03 Belgium 1997 - 2011 (10) 0.89 1.01 Bolivia 1998 - 2013 (13) 1.00 1.03 Botswana 1997 - 2000 (4) 0.75 1.10 Brazil 1998 - 2017 (16) 0.86 1.02 British Virgin Islands 1997 - 1999 (3) 0.11 1.62 Bulgaria 1997 - 2017 (15) 0.99 1.03 Burkina Faso 2001 - 2018 (14) 0.98 0.98 Cabo Verde 2009 - 2016 (8) 0.52 0.95 Cameroon 1998 - 2003 (5) 1.11 0.92 Canada 1974 - 2020 (37) 0.89 0.97 Chile 1998 - 2018 (16) 0.88 1.02 China 1993 - 2014 (20) 0.97 1.03 Colombia 1997 - 2018 (19) 0.91 1.10 Costa Rica 2014 - 2017 (4) 0.97 1.04 Cote d’Ivoire 1997 - 2000 (4) 0.98 1.59 Croatia 1997 - 2019 (18) 0.90 1.05 Curacao 2002 - 2016 (12) 0.94 1.11 Cyprus 1998 - 2019 (18) 0.91 1.05 Czech Republic 1995 - 2015 (14) 1.04 1.04 Denmark 1997 - 2020 (21) 0.91 0.97 Dominican Republic 2007 - 2016 (9) 0.77 1.10 Ecuador 2009 - 2018 (9) 0.66 0.84 Egypt 1998 - 2015 (17) 0.80 0.95 Estonia 1997 - 2019 (17) 0.96 1.02 Finland 1998 - 2020 (18) 0.97 1.03 France 1952 - 2018 (38) 0.87 1.02 Germany 1972 - 2020 (26) 0.89 1.01 Greece 1996 - 2019 (19) 0.90 1.04 Guatemala 2003 - 2019 (9) 0.91 1.02 Guinea 2005 - 2010 (5) 0.92 1.13 Honduras 2003 - 2015 (13) 0.80 0.97 Hong Kong 1997 - 2020 (21) 0.91 1.03 Hungary 1997 - 2019 (18) 0.94 1.06 Iceland 2003 - 2014 (12) 0.87 1.02 India 2000 - 2017 (13) 0.84 0.98 Iran 1997 - 2018 (21) 0.14 0.98 Ireland 1997 - 2019 (19) 0.94 0.99 Israel 1997 - 2018 (19) 0.98 1.17 Country Period $\overline{\varphi_{s,i}}$ $\overline{\varphi_{c,i}}$ Italy 1981 - 2020 (28) 0.93 1.01 Japan 1981 - 2018 (29) 0.97 1.03 Kazakhstan 1997 - 2018 (19) 0.97 1.03 Korea 1997 - 2018 (14) 0.97 1.01 Kuwait 2005 - 2017 (11) 1.07 1.07 Kyrgyzstan 1998 - 2019 (18) 0.72 1.06 Latvia 1998 - 2015 (16) 0.99 1.01 Lithuania 1997 - 2019 (16) 0.87 0.99 Luxembourg 1997 - 2018 (21) 0.97 1.01 Malaysia 2007 - 2015 (8) 0.80 1.10 Malta 1997 - 2019 (21) 0.77 0.99 Mexico 1997 - 2019 (21) 0.93 0.99 Moldova 1997 - 2018 (20) 0.80 0.93 Mongolia 2007 - 2019 (11) 0.95 1.02 Morocco 2000 - 2019 (14) 0.87 0.97 Netherlands 1997 - 2019 (16) 0.94 1.01 New Zealand 1997 - 2019 (17) 0.85 1.02 Nicaragua 2007 - 2018 (12) 0.90 0.99 Niger 1997 - 2019 (21) 0.87 0.96 Norway 1983 - 2020 (30) 0.82 1.01 Peru 2009 - 2019 (9) 0.96 1.02 Philippines 1997 - 2019 (18) 0.57 0.96 Poland 1997 - 2015 (13) 0.88 1.07 Portugal 1997 - 2020 (18) 0.98 1.02 Qatar 2004 - 2018 (12) 0.72 0.93 Romania 1997 - 2019 (20) 1.06 1.06 Russian Federation 1997 - 2017 (12) 0.99 1.06 Saudi Arabia 2004 - 2008 (3) 1.09 1.17 Senegal 2015 - 2015 (1) 0.03 0.78 Serbia 1999 - 2019 (18) 0.86 0.80 Slovakia 1997 - 2020 (15) 0.88 1.01 Slovenia 1997 - 2019 (17) 0.91 1.02 South Africa 1997 - 2020 (17) 0.96 1.03 Spain 1996 - 2017 (21) 0.95 1.04 Sweden 1951 - 2020 (65) 0.96 1.07 Switzerland 1997 - 2019 (20) 0.99 1.00 Tunisia 1997 - 2011 (12) 0.67 1.00 Turkey 2010 - 2017 (7) 0.64 1.02 USA 1972 - 2018 (37) 0.99 1.02 Ukraine 1997 - 2019 (23) 1.04 1.18 United Kingdom 1971 - 2018 (37) 0.99 1.02 Uzbekistan 2017 - 2017 (1) -0.01 0.67 Vanuatu 2003 - 2007 (4) 0.72 0.93 Venezuela 1998 - 2019 (20) 0.85 1.04
@@
Fig. 2: Thrift and free growth indexes derived from consumption (88 countries).
Table 3: Regression of $\Delta g(K)$ to $\theta_{c}$ and $\varphi_{c}$ (Screen = 0.01). $H_{0}\ \text{per thrift theory:}\ \theta_{c}=1\ \&\ \varphi_{c}=0$
$\theta_{c}$ $\varphi_{c}$
Regression of $\Delta g(K)$ to value shown ${}^{***}$ ${}^{***}$

### v4

@@ 1 Introduction and overview
Some other economists, including John Rae (1834) and John Stuart Mill (1848) (see Appendix C), argued that capital growth might also be explained by a rise in productivity of capital and labor already extant. Ways might found for existing factors to produce more, that is, and so to allow more consumption, or more capital growth, or any mix of the two, without inputs of net saving. Robert Solow (1957) allowed that possibility for “disembodied” growth, where plant and products already existing are repurposed or redeployed in more productive ways.
We test between those two explanations of capital growth, by net saving or by increase in productivity of capital and labor already in existence, by comparing net saving to concurrent change in market-value capital in 86 countries. As changes in net saving are expected to be associated with opposite changes in consumption, we also compare change in consumption to concurrent change in capital growth (capital acceleration). All data are drawn from national accounts of those countries as collated on the free website World Inequality Database.
@@
If free growth theory is right, tax policy and other policy to encourage saving over consumption should be reviewed. These policies include the higher tax on ordinary income than on capital gains, and the double tax on corporate dividends.
@@
2 Net saving, capital growth and capital acceleration
@@
$\Delta K_{i}=S_{net,i}+Q_{s,i}\ .$ (1)
Here $Q$ means the sum of market noise, which may prove positive or negative or zero, plus any part of capital growth explained by concurrent productivity gain as described by Rae and Mill, and by Solow as to disembodied growth. We call this sum of noise and productivity gain “free growth” in that it costs no net saving. $Q_{s,i}$ means its value as measured from data for saving in the $i^{th}$ year.
The object of testing is to find effects of $S_{net}$ on concurrent $\Delta K$, and so to help evaluate historic and current teachings as to those effects. We submitted above that most teaching, with exceptions noted as to Rae, Mill, and Solow, and within the warranted rate, predicts net saving to differ from concurrent growth in market-value capital only by market noise which tends to balance out over scale and time. Free growth $Q$ in Eq. (1), in that case, will give the market noise converging to zero. That consensus prediction, which we will challenge, will here be called “thrift theory”; $S_{net}$, in thrift theory, is expected to converge to $\Delta K$ if held within the warranted rate. That is,
$E(\Delta K)=S_{net}\quad\text{or equivalently}\quad E(Q_{s})=0,\quad\text{if}\quad\frac{\Sigma S_{net}}{\Sigma K}\leq u\ ,\quad\text{in thrift theory,}$, (2)
@@
$E(\Delta K)$ and $E(Q_{s})$ here give the expected values of $\Delta K$ and $Q_{s}$ respectively. Expected value means predicted average of outcomes over all observations. We will test those predictions against data for saving and capital growth taken from national accounts. As secular economic growth has tended to make later stocks and flows larger than earlier ones, we first divide by $K$ (normalize) to avoid overweighting of more recent years in finding that average. Division of Eq. (1) by $K_{i-1}$ gives
$\frac{{\Delta K}_{i}}{K_{i-1}}=\frac{S_{net,i}}{K_{i-1}}+\frac{Q_{s,i}}{K_{i-1}}\ .$ (3)
The first term in Eq. (3) gives capital growth rate $g(K)$. The second term is a variant of the Keynesian net saving rate $s_{net}$ where capital rather than output becomes the denominator. This flow will show as $s^{*}$, with the subscript “net” left implicit, and with the understanding that the denominator shows market-value capital rather than output. The third term in Eq. (3) will be called free growth rate and shown as $q_{s}$. Then
$g{(K)}_{i}=\frac{\Delta K_{i}}{K_{i-1}}\ ,\quad s^{*}_{i}=\frac{S_{net,i}}{K_{i-1}}\ ,\quad\text{and}\quad q_{i}=\frac{Q_{s,i}}{K_{i-1}}\ ,$
@@
$g(K)_{i}=s^{*}_{i}+q_{s,i}\ .$ (4)
By the definition $q_{s}=\frac{Q_{s}}{K}$, an expected value $E(q_{s})=0$ implies $E\left(q_{s}\right)=0$. Application of Eq. (2) to Eq. (4) now gives
$E\left(g(K)\right)=s^{*}\quad\text{and}\quad E(q_{s})=0\ ,\quad\text{under thrift assumptions,}$, (5)
where “thrift assumptions” are that thrift theory is correct and that the warranted rate is not exceeded. We will test Eq. (5) directly in Section 5 below. Meanwhile we will test it indirectly by testing its implications for changes in capital growth, which we call capital acceleration and show as $\Delta g(K)$. First note that Eq. (4) allows
$\Delta g{(K)}_{i}=\Delta s^{*}_{i}+\Delta q_{s,i}\ ,\quad\text{where}$ (6)
$\Delta g{(K)}_{i}=g{(K)}_{i}-g(K)_{i-1}\ ,\quad\Delta s^{*}_{i}=s^{*}_{i}-s^{*}_{i-1}\ ,\quad\text{and}\quad\Delta q_{s,i}=q_{s,i}-q_{s,i-1}\ .$
@@
$E\left(\Delta g(K)\right)=\Delta s^{*}\quad\text{and}\quad E(\Delta q_{s})=0\ ,\quad\text{under thrift assumptions.}$. (7)
Division of Eq. (6) by capital acceleration, and rearrangement, gives
$\frac{\Delta s^{*}_{i}}{\Delta g{(K)}_{i}}+\frac{\Delta q_{s,i}}{\Delta g{(K)}_{i}}=1\ .$ (8)
Define $\theta_{s,i}=\frac{\Delta s^{*}_{i}}{\Delta g{(K)}_{i}}$ and $\varphi_{s,i}=\frac{\Delta q_{s,i}}{\Delta g{(K)}_{i}}$ to restate Eq. (8) as
@@
$\theta$ and $\varphi$ will be called the “thrift index” and “free growth index” respectively. $\theta_{s}$ and $\varphi_{s}$ give their values as found from data for change in net saving. Expected values, again, are predicted averages of outcomes. Thus Eq. (11) and thrift theory can be tested by finding average values of $\theta_{s,i}$ and comparing findings to the expected value $\theta=1$. First we find
@@
$\overline{\theta_{s,i}}\cong 1\quad\text{and}\quad\overline{\varphi_{s,i}}\cong 0\ ,\quad\text{under thrift assumptions.}$.
@@
Fig. 1 shows average values of $\theta_{s,i}$ and $\varphi_{s,i}$ for 86 countries, both unweighted and weighted to GDP, over the period 1980-2022. To control distortions brought by small absolute denominators, years were screened out where $|\Delta g(K)|$ was found at less than 0.01 (see Section 10). Results show $\overline{\varphi_{s,i}}\cong 1$ and $\overline{\theta_{s,i}}\cong 0$. These findings appear to refute thrift theory, and to support free growth theory as defined earlier. $Q_{s}$, predicted in thrift theory to describe effects of market noise converging to zero, is revealed to include also the effects of productivity gain as described by Rae, Mill and Solow. We will now test thrift theory from a different approach.
Fig. 1: Thrift and free growth indexes derived from net saving (86 countries).
Table 1: Regression of $\Delta s^{*}$ and $\Delta q_{s}$ on $\Delta g(K)$ (Screen = 0.01). $H_{0}\ \text{per thrift theory:}\ \Delta g(K)\cong\Delta s^{*}\ \&\ \Delta q_{s}\cong 0$
$\Delta s^{*}$ $\Delta q_{s}$
Regression of value shown on $\Delta g(K)$ 0.1092∗∗∗ 0.8908∗∗∗
@@ 3 Consumption and capital growth
In national accounts and most of the literature, saving equals consumption forgone in that a fixed source of spending is spent on the sum of the two. Replacement saving to offset depreciation costs an equal potential gain in consumption while making up for losses in capital, that is, and net saving costs an equal actual decrease in consumption. We accept these meanings as definitions of saving, while disputing only the doctrine that investment of net saving brings actual increase in capital. Keynes (1936) expressed this definition in the equation $c+s=1$, where $c$ and $s$ are the ratios of consumption and gross saving to gross output, or consumption and net saving to net output. We have been dividing by capital, rather than output, and will let $c^{*}$ show the ratio of consumption to capital. We then adapt the Keynesian equation to show
$s^{*}_{i}+c^{*}_{i}=1\ ,\quad\text{from which}\ s^{*}_{i}=1-c^{*}_{i}\ ,\quad\text{where}\ c^{*}_{i}=\frac{C_{i}}{K_{i-1}}\ .$
We now define $q_{c}$ as free growth rate $q$ found from data for consumption rather than from saving. By analogy to Eqs. (4) and (5), then,
$g(K)_{i}=1-c_{i}^{*}+q_{c,i}\ ,\quad\text{and}$ (12)
$E(g(K))=1-c^{*}\quad\text{and}\quad E(q_{c})=0\quad\text{under thrift assumptions.}$. (13)
Meanwhile Eq. (12) allows
@@
$E(\Delta g(K))=-\Delta c^{*}\quad\text{and}\quad E(\Delta q_{c})=0\quad\text{under thrift assumptions.}$. (15)
As with Eqs. (5) and (7), we will put off direct tests of Eq. (13) until Section 5, and meanwhile test it indirectly by testing its implications from Eq. (14) forward. As with Eq. (6), we first divide Eq. (14) by capital acceleration $\Delta g(K)$ to find
$\frac{\Delta q_{c}}{\Delta g(K)_{i}}+\frac{-\Delta c^{*}_{i}}{\Delta g(K)_{i}}=1\ .$ (16)
Define $\theta_{c,i}=\frac{-\Delta c^{*}_{i}}{\Delta g(K)_{i}}$ and $\varphi_{c,i}=\frac{\Delta q_{c,i}}{\Delta g(K)_{i}}$ to re-express Eq. (16) as
@@
Table 2: Average $\varphi_{s,i}$ and $\varphi_{c,i}$ in 86 countries (screen = 0.01). Number of years clearing screen shown in ()
@@
Country Period $\overline{\varphi_{s,i}}$ $\overline{\varphi_{c,i}}$ Armenia 1997 - 2018 (17) 0.94 1.30 Aruba 1997 - 2001 (5) -0.38 8.98 Australia 1962 - 2019 (41) 0.97 1.09 Austria 1997 - 2013 (11) 0.93 1.11 Azerbaijan 1997 - 2018 (22) 0.39 1.85 Bahrain 2010 - 2013 (4) 0.59 1.32 Belgium 1997 - 2014 (11) 0.90 1.06 Bolivia 1999 - 2013 (12) 1.03 1.28 Botswana 1997 - 2000 (4) 0.73 1.27 Brazil 1998 - 2018 (20) 0.86 1.21 British Virgin Islands 1997 - 1999 (3) 0.35 3.63 Bulgaria 1997 - 2017 (15) 0.99 1.21 Burkina Faso 2001 - 2018 (15) 0.94 1.23 Cameroon 1998 - 2003 (6) 0.35 4.59 Canada 1974 - 2020 (37) 0.91 1.11 Cape Verde 2009 - 2017 (9) 0.56 2.01 Chile 1998 - 2018 (18) 0.93 1.37 China 1993 - 2014 (19) 0.91 1.04 Colombia 1997 - 2019 (17) 0.98 2.26 Costa Rica 2014 - 2017 (4) 0.97 1.09 Croatia 1997 - 2018 (21) 0.83 2.09 Curaçao 2002 - 2016 (12) 1.13 1.18 Cyprus 1997 - 2018 (19) 0.94 1.11 Czechia 1995 - 2018 (14) 0.99 1.04 Côte d’Ivoire 1997 - 2000 (4) 1.03 1.10 Denmark 1999 - 2020 (20) 0.92 1.00 Dominican Republic 2007 - 2015 (8) 0.76 1.33 Ecuador 2009 - 2018 (10) 0.85 1.87 Egypt 1998 - 2015 (17) 1.04 1.75 Estonia 1997 - 2017 (15) 0.93 1.11 Finland 1997 - 2020 (19) 0.97 1.11 France 1952 - 2019 (41) 0.87 1.13 Germany 1971 - 2013 (22) 1.01 1.21 Greece 1997 - 2019 (21) 0.90 1.27 Guatemala 2007 - 2019 (9) 0.96 0.93 Guinea 2005 - 2010 (6) 0.94 1.27 Honduras 2002 - 2015 (14) 0.70 1.56 Hong Kong 1997 - 2020 (21) 0.91 1.12 Hungary 1997 - 2018 (17) 0.93 1.11 Iceland 2002 - 2014 (13) 0.85 1.07 India 2000 - 2017 (15) 0.88 1.08 Iran 1998 - 2018 (19) 0.15 1.09 Ireland 1997 - 2019 (20) 0.96 0.96 Country Period $\overline{\varphi_{s,i}}$ $\overline{\varphi_{c,i}}$ Israel 1997 - 2017 (18) 0.95 1.68 Italy 1982 - 2017 (21) 1.01 0.98 Japan 1981 - 2017 (28) 0.95 1.13 Kazakhstan 1997 - 2018 (18) 0.90 1.09 Kuwait 2005 - 2017 (11) 0.98 1.12 Kyrgyzstan 1998 - 2019 (20) 0.73 1.19 Latvia 1998 - 2015 (16) 1.00 1.13 Lithuania 1997 - 2016 (16) 0.93 1.14 Luxembourg 1997 - 2018 (18) 0.99 1.09 Malaysia 2008 - 2015 (7) 1.04 1.27 Malta 1997 - 2019 (21) 0.80 1.03 Mexico 1997 - 2019 (21) 0.95 1.14 Moldova 1997 - 2018 (21) 0.74 1.19 Mongolia 2008 - 2019 (10) 0.92 1.07 Morocco 2000 - 2019 (16) 0.84 1.38 Netherlands 1997 - 2019 (14) 0.94 1.12 New Zealand 1997 - 2019 (21) 0.92 1.12 Nicaragua 2008 - 2018 (11) 0.93 1.13 Niger 1997 - 2019 (20) 0.82 1.95 Norway 1983 - 2020 (32) 0.86 1.12 Peru 2009 - 2019 (9) 0.96 1.33 Philippines 1997 - 2019 (19) 0.68 2.20 Poland 1997 - 2019 (15) 0.91 1.23 Portugal 1997 - 2020 (17) 0.98 1.08 Qatar 2004 - 2018 (12) 0.72 0.87 Romania 1997 - 2019 (16) 1.02 1.11 Russia 1997 - 2018 (14) 0.89 1.11 Saudi Arabia 2005 - 2009 (3) -0.92 0.21 Serbia 1999 - 2019 (16) 0.59 2.04 Slovakia 1997 - 2020 (19) 0.93 1.20 Slovenia 1997 - 2019 (17) 0.92 1.08 South Africa 1997 - 2019 (17) 0.95 1.14 South Korea 1997 - 2018 (14) 0.96 1.04 Spain 1997 - 2017 (20) 0.95 1.15 Sweden 1951 - 2020 (57) 0.98 1.05 Switzerland 1993 - 2018 (22) 0.94 1.03 Tunisia 1997 - 2011 (12) 0.69 1.06 Turkey 2011 - 2017 (7) 0.83 1.24 USA 1972 - 2018 (40) 1.00 1.14 Ukraine 1997 - 2019 (22) 0.93 1.16 United Kingdom 1971 - 2018 (38) 1.00 1.14 Vanuatu 2003 - 2007 (4) 0.72 0.92 Venezuela 1999 - 2019 (19) 0.05 1.61
@@
Fig. 2: Thrift and free growth indexes derived from consumption (86 countries).
Table 3: Regression of $-\Delta c^{*}$ and $\Delta q_{c}$ on $\Delta g(K)$ (Screen = 0.01). $H_{0}\text{ per thrift theory: }\Delta g(K)\cong-\Delta c^{*}\&\Delta q_{c}\cong 0$
$-\Delta c^{*}$ $\Delta q_{c}$
Regression of value shown on $\Delta g(K)$0.4006∗∗∗ 1.401∗∗∗